OpenStudy (anonymous):

help please ! Find the 9th term of the arithmetic sequence. a10= 100 and a20= 50; a9=??? formula is an=a1+(n-1)d i found the common difference where d=-5 i just need help finding a1 ...

OpenStudy (cwrw238):

d = -5 thats correct a10 = a1 + 9d = 100 so a1 + 9(-5) = 100 solve this for a1

OpenStudy (anonymous):

what i did was 100=a1+(10-1)(-5) 100=a1+(-45)

OpenStudy (cwrw238):

that right as far as it goes

OpenStudy (anonymous):

so do i distribute the + to -45 ?

OpenStudy (cwrw238):

100 = a1 - 45 now add 45 to both sides

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

145 = a1

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

okay cool. but can you please help me go on and find the 9th term ?

OpenStudy (cwrw238):

nth term = a1 + (n-1)d 9th term = 145 + (9 - 1)* -5

OpenStudy (anonymous):

so a9 = 105

OpenStudy (anonymous):

thank you!

OpenStudy (cwrw238):

right - yw